Circular No. 6634

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) G. Cremonese, Astronomical Observatory, Padua; H. Rauer, Observa-

toire de Paris, Meudon; and A. Fitzsimmons, Queen’s University, Belfast, on behalf of the European Hale-Bopp Team, report further observations pertaining to the neutral-sodium tail recently discovered in C/1995 O1

(IAUC 6631; cf. http://ing.iac.es/PR/Caption ING397.html): “Spec-

tra were obtained on Apr. 19.90 UT with the Utrecht Echelle Spectrograph at the 4.2-m William Herschel Telescope at the Observatorio del Roque de los Muchachos, La Palma, with a 4′′.0 × 1′′.1 slit positioned 1^{o}.5 along the sodium tail at p.a. 56^{o} (this position angle being determined by simultaneous imaging with the CoCAM instrument). The location of the slit corresponds to a cometocentric distance of 1.1 × 10^{7} km along the tail. A series of spectral windows were observed between 500 nm and 800 nm at a resolution of 7 km/s. In a 15-min exposure, strong Na D lines were detected (their strength relative to the sky lines being ∼ 1.3 : 1), redshifted at a line-of-sight velocity of +105 km/s, which corresponds to a cometocentric velocity of ∼ 95 km/s along the tail. The lines are resolved with an intrinsic FWHM of ∼ 11 km/s. These data confirmed that neutral Na is present in the newly-detected third type of cometary tail. Emissions from H_{2}O^{+}, NH_{2} and C_{2} (among other molecules) were not detected at this tail position. The same instrumentation was used on Apr. 20.85 UT to obtain spectra at 10′′ from the nucleus and at 0^{o}.7 (cometocentric distance 5.0 × 10^{6} km) along the Na tail. Once again, Na D-line emission associated with the comet was observed, at cometocentric velocities of 0 and 58 km/s, respectively. From the measured velocity distribution and surface brightness along the tail, it would appear that the data are consistent with a model in which neutral sodium atoms are released from still-unidentified sources within the coma and then accelerated in the antisolar direction by simple fluorescence. At the same time as the Apr. 19.90 spectroscopic data were being secured, CCD imaging was performed by J. Licandro, Instituto de Astrofisíca de Canarias, with the 1-m Jacobus Kapteyn Telescope through a narrow-band filter (λ_{0} = 587.7 nm, FWHM = 4.9 nm) and an adjacent continuum filter. A comparison between these images over the 5′ field of view did not reveal any features that could be identified as peculiar to neutral Na.”

COMET 29P/SCHWASSMANN-WACHMANN 1 Visual m_{1} estimates: Mar. 10.93 UT, 13.5 (K. Hornoch, Lelekovice,

Czech Republic, 0.35-m refl.); 28.93, 14.2 (J. Carvajal, Teruel, Spain, 0.44m refl.); Apr. 1.92, [13.5 (K. Sárneczky, Ráktanya, Hungary, 0.44-m refl.).

1997 April 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden